To practice Problem-Solving Strategy 3.1 Projectile motion problems.

A ball thrown horizontally with speed viviv_i = 20.0 m/sm/s travels a horizontal distance of ddd = 58.0 mm before hitting the ground. From what height hhh was the ball thrown?

If it takes a ball dropped from rest 2.900 s to fall to the ground, from what height H was it released? Express your answer in meters.
Answer #1

horizontal velocity = Vix = 20 m/sec

initial vertical velocity = Viy = 0 m/sec

horizontal distance = x = 58 m

horizontal direction acceleration ax = 0

                         x = Vix * t = 58 = 20 * t

                          t = 58 /20 = 2.9 sec

height = h = Viy *t + 1/2 g t^2

             h = 0 + 1/2 *9.8* 2.9^2 = 41.2 = 41 m

b) V0 = 0 ( REST)

    t = 2.9 sec

    h = v0 * t + 1/2 g t^2

   h = 0+ 1/2 *9.8*2.9^2 = 41.2 m = 41 m